# MAViS: Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays

### v1

@@ Abstract
Arrays of gate-defined semiconductor quantum dots are among the leading candidates for building scalable quantum processors. High-fidelity initialization, control, and readout of spin qubit registers require exquisite and targeted control over key Hamiltonian parameters that define the electrostatic environment. However, due to the tight gate pitch, capacitive crosstalk between gates hinders independent tuning of chemical potentials and interdot couplings. While virtual gates offer a practical solution, determining all the required cross-capacitance matrices accurately and efficiently in large quantum dot registers is an open challenge. Here, we establish a Modular Automated Virtualization System (MAViS) – a general and modular framework for autonomously constructing a complete stack of multi-layer virtual gates in real time. Our method employs machine learning techniques to rapidly extract features from two-dimensional charge stability diagrams. We then utilize computer vision and regression models to self-consistently determine all relative capacitive couplings necessary for virtualizing plunger and barrier gates in both low- and high-tunnel-coupling regimes. Using MAViS, we successfully demonstrate accurate virtualization of a dense two-dimensional array comprising ten quantum dots defined in a high-quality Ge/SiGe heterostructure. Our work offers an elegant and practical solution for the efficient control of large-scale semiconductor quantum dot systems.
@@ I Introduction
Qubits that utilize the spin degree of freedom of charge carriers confined within semiconductor quantum dots (QDs) show great promise for practical large-scale quantum computation [1]. In the last decade, silicon spin qubit systems based on one-dimensional arrays have demonstrated long coherence times, high-fidelity single- and two-qubit gates, coherent quantum information transfer and compatibility with industrial manufacturing techniques [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. In more recent years, planar germanium QD hole-spin qubits have emerged as an alternative semiconductor technology that can ease certain challenges in qubit control and device engineering [13]. In particular, the strong spin-orbit coupling and small effective mass of holes in germanium [14, 15], together with a highly uniform and low noise material platform [16, 17, 18], have sparked tunable QDs [19, 20, 21] and spin qubits systems arranged in two dimensions [22, 23, 24]. As compelling fabrication methods further advance providing routes for engineering of large QDs registers [25, 26], it is expected that automation will be playing a pivotal role in orchestrating precise quantum operations at various levels [27, 28, 29, 30, 31, 32, 33, 34, 35, 11].
@@
Table 1: Table illustrating the layers of our virtualization approach.
Virt. layer Description Output gates Input gates Trans.
@@
In this work, we advance this effort by combining state-of-the-art QD arrays with modern machine learning to address the open challenge of virtualization. We propose and validate a modular automated virtualization system (MAViS), a framework to achieve orthogonal control of the electrostatic potential landscape. We test-bed the approach on a two-dimensional (2D) germanium ten QD array by defining a multi-layer stack of virtual gates that accurately control the QD energies and couplings. Here, we take advantage of prior works aiming to automate the operation of semiconductor qubits to build a framework that can efficiently fine-tune multi-QD devices.
Initial virtualization methods relied on laboratory heuristics, along with device-specific information processing [43, 44]. With the advent of modern machine learning (ML), automated methods relying on ML models and traditional curve fitting to identify orthogonal plunger gates and mitigate capacitative couplings were first proposed in Refs. [45, 46], improving on minimal computer-automated tuning algorithms [47, 48]. The prescribed methods extract the inclinations of transition lines from 2D charge stability diagrams (CSDs) to define the virtualization matrix. However, as QDs in large arrays decouple from the reservoirs, causing a reduced exchange rate of carriers with respect to the typical measurement scan rates, the 2D CSDs become more challenging to interpret and analyze due to, e.g., secondary charge transitions and more pronounced latching transition lines [20]. Our method is articulated on several virtualization layers, exploits both ML and classic analysis techniques for robustness, and has a larger scope than existing approaches, extending beyond plunger gate virtualization. By ultimately enabling the definition of virtual barrier gates that tune the interdot couplings over large ranges, without affecting the QD potentials, we build an operation recipe that provides control over the key Hamiltonian parameters defining the electrostatics of the QD system. Finally, our ability to accurately decode information from complex 2D CSDs enables us to study the boundaries of the linear virtual gates approach and provide a method for using control parameters beyond such limits. Our work demonstrates the full virtualization of a dense 2D array of ten QDs in a planar germanium quantum well in the few-hole regime. We start with a device hosting weakly-coupled QDs pre-tuned via manual operations and a set of unvirtualized sensor, plunger, and barrier gates. By leveraging the MAViS, we then seamlessly transition into a fully orthogonal gate space, achieving complete virtualization in an automated and efficient manner.
The organization of the paper is as follows: In Sec. II, we present an overview of the virtualization framework, including the description of the full virtualization stack in Sec. II.1, the experimental setup used to test the MAViS in Sec. II.2, and the data processing techniques in Sec. II.3. The plunger- and barrier-specific tools are described in Sec. II.4 and Sec. II.5, respectively. The performance of the MAViS in autonomously defining a set of the virtual plunger and barrier gates to control the ten-QD device is presented in Sec. III. Finally, in Sec. IV we summarize our results and discuss future outlook.
II Methods
A typical multi-QD spin qubit device is controlled by a set of $N_{\mathrm{P}}$ plunger gates $\mathbf{P}$, $N_{\mathrm{B}}$ barrier gates $\mathbf{B}$, and $N_{\mathrm{S}}$ charge-sensing plunger gates $\mathbf{S}$. Tuning the device into a QD regime requires an extensive search over relatively large voltage ranges, with the capacitive crosstalk between gates further complicating the search. We propose a multi-layer virtualization stack that is both device-agnostic and modular, as illustrated in Table 1.
Figure 1: Device architecture and virtualization workflow. (a) Layout of a 10 QD array based on a 3-4-3 geometry. Holes are trapped in gate-defined germanium QDs controlled by a set of barrier (dark blue), plunger (magenta), and screening (cyan) gates. (b) Schematic displaying the approximate potential landscape and position of the QDs in the array, with Dn, for $n\in[1,\dots,10]$, indicating each dot and CSN, CSE, CSS, and CSW indicating the north, east, south, and west charge sensor, respectively. (c) A typical 2D CSD of a double QD measured via the rf-reflectometry charge sensing on CSN. (d) The workflow of the ML model: each pixel is assigned a probability to be part of a horizontal, vertical, and interdot transition (diagonal and no transition classes not shown). (e) Probability distribution of each pixel to be a horizontal (left) or vertical (right) transition class. (f) Probability distribution of each pixel to be an interdot class. (g) A Gaussian fit to the log-transformed probability distribution for the interdot class. (h) Extracted coordinates of the interdot centers of mass based on dynamic thresholding.
@@ II.1 Virtualization stack
To ensure the high sensitivity of the charge sensors while the QDs are calibrated, it is convenient to first address the compensation to the charge sensor plunger gates by exploiting a set of virtual gates $\{\mathbf{vP},\mathbf{vB}\}=\{\mathrm{vP}_{i},\mathrm{vB}_{j}\,|\,i=1,\dots,N_{\mathrm{P}};\,j=1,\dots,N_{\mathrm{B}}\}$, that can be used to maintain the charge sensors tuned to their most sensitive voltage points (layer 1 in Table 1). Achieving this step requires knowledge of the relative lever arms from the plunger and barrier gates controlling the QDs to the plunger gates of the charge sensors ($\mathrm{S}_{k}$, with $k\in[1,\cdots,N_{\mathrm{S}}]$). The link between sensor-virtualized gates and real gates is defined through the virtual matrix $M_{1}$ and reads:
@@
with the absolute value of the entries of the inverse matrix $M^{-1}_{1}$ being the relative lever arms of the QD plunger gates ($\alpha_{k,N_{\mathrm{S}}+i}^{S}$) and of the barrier gates ($\alpha_{k,N_{S}+N_{\mathrm{P}}+j}^{S}$), that are extracted experimentally. As an example, if an increase of $1$ $\mathrm{mV}$ in the barrier $B_{1}$ produces a shift of $-0.2$ $\mathrm{mV}$ to the position of Coulomb peak associated with the third charge sensor in the $S_{3}$ space, the entry $(3,N_{S}+N_{\mathrm{P}}+1)$ of $M^{-1}_{1}$ reads $-0.2$, and $0.2$ is the relative lever arm of $B_{1}$ on $S_{3}$.
We also note that splitting the virtualization stack into multiple layers is advantageous for keeping virtual matrices in basic forms. While $M^{-1}_{1}$ appears as a large matrix, in practice, it is a diagonal matrix exhibiting off-diagonal elements in only the first $N_{\mathrm{S}}$ rows. Furthermore, any control parameter built as a linear combination of the $\mathbf{vP}$ and $\mathbf{vB}$ virtual gates will also keep all charge sensors tuned. This allows building up higher-hierarchy virtualization layers using the newly defined $\{\mathbf{vP},\mathbf{vB}\}$ gates without further adjusting the compensations to the charge sensors.
For spin control and readout, it is practical to operate in an orthogonal plunger gate space where each plunger controls only the designated site’s chemical potential independently from the other gates. This is addressed in layer 2 of MAViS, which we call plunger gate orthogonalization, where a set of ad-hoc virtual plunger gates $\mathbf{O}$ provide orthogonal control over the chemical potentials through a second virtual matrix $M_{2}$ as:
@@
with the absolute value of the entry of the inverse matrix $M^{-1}_{2}$, $\alpha^{O}_{n,i}$, being the relative lever arm of $\mathrm{vP}_{i}$ to dot $n$.
@@
with the entries $\alpha^{N}_{i,i}$ of of the inverse matrix $M^{-1}_{3}$ defined as $d_{i}/V_{D}$ with $V_{D}$ the target charging voltage, and $d_{i}$ the charging voltage measured along the $\mathrm{O}_{i}$ voltage space.
@@
where the entry of $M^{-1}_{4}$ at the position $\alpha_{N_{\mathrm{B}}+i,j}^{OFF}$ encodes the relative shift of the charge state of dot $i$ as the barrier $\mathrm{vB}_{j}$ voltage is changed.
The execution of fast two-qubit exchange-based gates requires modulating the tunnel coupling between adjacent QDs across several orders of magnitudes [36, 22, 1]. A large voltage pulse on the barrier gates also affects the effective QD positions as the wave functions are brought close to overlap. This effect is expected to modify the aforementioned capacitive couplings limiting the range of validity of the virtual matrix extracted in the OFF-coupling regime. As a consequence, barrier control compensations dedicated to the high-coupling (ON) regime are introduced in layer 5. We anticipate that, depending on the targeted exchange interactions, linear compensations may not suffice and that non-linear corrections to the plunger gates may be necessary when a barrier is pulsed substantially. Therefore, we heuristically define quadratically-corrected virtual gates $\mathbf{K}$ as follows:
$\mathrm{K}_{j}=\mathrm{J}_{j}+\sum_{m}\left[\alpha^{ON}_{j,m}{\mathrm{N}_{m}}^{2}+\beta^{ON}_{j,m}\mathrm{N}_{m}\right],\vspace{-3pt}$ (5)
where the $(\sum_{m})$ involves only the nearest-neighboring plunger gates to barrier gate $j$, and $\alpha^{ON}_{j,m}$, $\beta^{ON}_{j,m}$ represent the first and second order compensation coefficients of barrier $J_{j}$ to the QDs $m$, respectively.
@@ II.2 Experimental setup
The full virtualization approach is demonstrated on a dense array of ten QDs defined in a planar germanium quantum well. This device [50] is based on low-disorder Ge/SiGe heterostructure grown on a Ge substrate, with the quantum well separated from the dielectric interface by a $55$ $\mathrm{nm}$ SiGe barrier [18].
The device uses a multilayer gate architecture to define an array of ten QDs arranged in a 3-4-3 configuration, see Fig. 1(a). Ten plunger gates and twelve barrier gates labeled $\mathrm{P}_{i}$ and $\mathrm{B}_{j}$ for $i\in[1,\dots,10]$ and $j\in[1,\dots,12]$, respectively, offer control of the array’s electrostatic potential landscape. Four additional plunger gates, labeled $\mathrm{P}_{N}$, $\mathrm{P}_{E}$, $\mathrm{P}_{W}$, and $\mathrm{P}_{S}$, control the sensor QD’s potentials. The ten QDs are labeled Dn, with $n\in[1,\dots,10]$, and their four peripheral sensor QDs are labeled CSN, CSE, CSW, and CSS, based on their cardinal directions, see Fig. 1(b). The additional eight screening gates screen the electric field from the plunger gates to prevent the formation of spurious QDs. The screening gates are omitted from the virtualization stack as they are not changed during normal device operation.
Prior to executing the virtualization stack, the QD array is pre-tuned to the few-hole regime, with each QD being either in a single-, triple-, or quintuple-hole occupancy. The four charge sensors permit fast simultaneous radio-frequency (rf)-reflectometry charge sensing in combination with video-mode acquisition and frequency multiplexing [51, 52, 53].

### v2

# Modular Autonomous Virtualization System for Two-Dimensional Semiconductor Quantum Dot Arrays

@@ Abstract
Arrays of gate-defined semiconductor quantum dots are among the leading candidates for building scalable quantum processors. High-fidelity initialization, control, and readout of spin qubit registers require exquisite and targeted control over key Hamiltonian parameters that define the electrostatic environment. However, due to the tight gate pitch, capacitive crosstalk between gates hinders independent tuning of chemical potentials and interdot couplings. While virtual gates offer a practical solution, determining all the required cross-capacitance matrices accurately and efficiently in large quantum dot registers is an open challenge. Here, we establish a modular automated virtualization system (MAViS)—a general and modular framework for autonomously constructing a complete stack of multilayer virtual gates in real time. Our method employs machine learning techniques to rapidly extract features from two-dimensional charge stability diagrams. We then utilize computer vision and regression models to self-consistently determine all relative capacitive couplings necessary for virtualizing plunger and barrier gates in both low- and high-tunnel-coupling regimes. Using MAViS, we successfully demonstrate accurate virtualization of a dense two-dimensional array comprising ten quantum dots defined in a high-quality Ge/SiGe heterostructure. Our work offers an elegant and practical solution for the efficient control of large-scale semiconductor quantum dot systems.
@@ I Introduction
Qubits that utilize the spin degree of freedom of charge carriers confined within semiconductor quantum dots (QDs) show great promise for practical large-scale quantum computation [1]. In the past decade, silicon spin qubit systems based on one-dimensional arrays have demonstrated long coherence times, high-fidelity single- and two-qubit gates, coherent quantum information transfer and compatibility with industrial manufacturing techniques [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. In more recent years, planar germanium QD hole-spin qubits have emerged as an alternative semiconductor technology that can ease certain challenges in qubit control and device engineering [13]. In particular, the strong spin-orbit coupling and small effective mass of holes in germanium [14, 15], together with a highly uniform and low noise material platform [16, 17, 18], have sparked tunable QDs [19, 20, 21] and spin qubits systems arranged in two dimensions [22, 23, 24]. As compelling fabrication methods further advance providing routes for engineering of large QD registers [25, 26], it is expected that automation will be playing a pivotal role in orchestrating precise quantum operations at various levels [27, 28, 29, 30, 31, 32, 33, 34, 35, 11].
@@
In this work, we advance this effort by combining state-of-the-art QD arrays with modern machine learning (ML) to address the open challenge of virtualization. We propose and validate a modular automated virtualization system (MAViS), a framework to achieve orthogonal control of the electrostatic potential landscape. We test bed the approach on a two-dimensional (2D) germanium ten QD array by defining a multilayer stack of virtual gates that accurately control the QD energies and couplings. Here, we take advantage of prior works aiming to automate the operation of semiconductor qubits to build a framework that can efficiently fine-tune multi-QD devices.
Initial virtualization methods relied on laboratory heuristics, along with device-specific information processing [43, 44]. With the advent of modern machine learning, automated methods relying on ML models and traditional curve fitting to identify orthogonal plunger gates and mitigate capacitative couplings were first proposed in Refs. [45, 46], improving on minimal computer-automated tuning algorithms [47, 48]. The prescribed methods extract the inclinations of transition lines from 2D charge stability diagrams (CSDs) to define the virtualization matrix. However, as QDs in large arrays decouple from the reservoirs, causing a reduced exchange rate of carriers with respect to the typical measurement scan rates, the 2D CSDs become more challenging to interpret and analyze due to, e.g., secondary charge transitions and more pronounced latching transition lines [20]. Our method is articulated on several virtualization layers, exploits both ML and classic analysis techniques for robustness, and has a larger scope than existing approaches, extending beyond plunger gate virtualization. By ultimately enabling the definition of virtual barrier gates that tune the interdot couplings over large ranges, without affecting the QD potentials, we build an operation recipe that provides control over the key Hamiltonian parameters defining the electrostatics of the QD system. Finally, our ability to accurately decode information from complex 2D CSDs enables us to study the boundaries of the linear virtual gates approach and provide a method for using control parameters beyond such limits. Our work demonstrates the full virtualization of a dense 2D array of ten QDs in a planar germanium quantum well in the few-hole regime. We start with a device hosting weakly coupled QDs pretuned via manual operations and a set of unvirtualized sensor, plunger, and barrier gates. By leveraging the MAViS, we then seamlessly transition into a fully orthogonal gate space, achieving complete virtualization in an automated and efficient manner.
The organization of the paper is as follows: In Sec. II, we present an overview of the virtualization framework, including the description of the full virtualization stack in Sec. II.1, the experimental setup used to test the MAViS in Sec. II.2, and the data processing techniques in Sec. II.3. The plunger- and barrier-specific tools are described in Secs. II.4 and II.5, respectively. The performance of the MAViS in autonomously defining a set of the virtual plunger and barrier gates to control the ten-QD device is presented in Sec. III. Finally, in Sec. IV we summarize our results and discuss future outlook.
II Methods
A typical multi-QD spin qubit device is controlled by a set of $N_{\mathrm{P}}$ plunger gates $\mathbf{P}$, $N_{\mathrm{B}}$ barrier gates $\mathbf{B}$, and $N_{\mathrm{S}}$ charge-sensing plunger gates $\mathbf{S}$. Tuning the device into a QD regime requires an extensive search over relatively large voltage ranges, with the capacitive crosstalk between gates further complicating the search. We propose a multilayer virtualization stack that is both device agnostic and modular, as illustrated in Table 1.
Table 1: The five virtualization layers implemented in MAViS. The first column indicates the layer number. The second column gives a description of each layer. Columns three and four give the output and input gates for each layer, respectively. The equation defining the transformation between the input and output gates is given in the last column.
Layer Description Output gates Input gates Trans.
@@ II.1 Virtualization stack
To ensure the high sensitivity of the charge sensors while the QDs are calibrated, it is convenient to first address the compensation to the charge sensor plunger gates by exploiting a set of virtual gates $\{\mathbf{vP},\mathbf{vB}\}=\{\mathrm{vP}_{i},\mathrm{vB}_{j}\,|\,i=1,\dots,N_{\mathrm{P}};\,j=1,\dots,N_{\mathrm{B}}\}$ that can be used to maintain the charge sensors tuned to their most sensitive voltage points (layer 1 in Table 1). Achieving this step requires knowledge of the relative lever arms from the plunger and barrier gates controlling the QDs to the plunger gates of the charge sensors ($\mathrm{S}_{k}$, with $k\in[1,\dots,N_{\mathrm{S}}]$). The link between sensor-virtualized gates and real gates is defined through the virtual matrix $M_{1}$ and reads:
@@
with the absolute value of the entries of the inverse matrix $M^{-1}_{1}$ being the relative lever arms of the QD plunger gates ($\alpha_{k,N_{\mathrm{S}}+i}^{S}$) and of the barrier gates ($\alpha_{k,N_{S}+N_{\mathrm{P}}+j}^{S}$) that are extracted experimentally. As an example, if an increase of $1$ $\mathrm{mV}$ in the barrier $\mathrm{B}_{1}$ produces a shift of $-0.2$ $\mathrm{mV}$ to the position of Coulomb peak associated with the third charge sensor in the $\mathrm{S}_{3}$ space, the entry $(3,N_{S}+N_{\mathrm{P}}+1)$ of $M^{-1}_{1}$ reads $-0.2$, and $0.2$ is the relative lever arm of $\mathrm{B}_{1}$ on $\mathrm{S}_{3}$.
We also note that splitting the virtualization stack into multiple layers is advantageous for keeping virtual matrices in basic forms. While $M^{-1}_{1}$ appears as a large matrix, in practice, it is a diagonal matrix exhibiting off-diagonal elements in only the first $N_{\mathrm{S}}$ rows. Furthermore, because each of the sensor-virtualized gates $\mathbf{vP}$ and $\mathbf{vB}$ can (by design) be varied without otherwise changing the compensations to the charge sensors, any control parameter built as a linear combination of these virtual gates will also keep all charge sensors tuned. This allows building up higher-hierarchy virtualization layers using the newly defined $\{\mathbf{vP},\mathbf{vB}\}$ gates, while the sensor compensations are adjusted automatically as part of the virtual gates.
For spin control and readout, it is practical to operate in an orthogonal plunger gate space where each plunger controls only the designated site’s chemical potential independently from the other gates. This is addressed in layer 2 of MAViS, which we call plunger gate orthogonalization, where a set of ad hoc virtual plunger gates $\mathbf{O}$ provide orthogonal control over the chemical potentials through a second virtual matrix $M_{2}$ as
@@
with the absolute value of the entry of the inverse matrix $M^{-1}_{2}$, $\alpha^{O}_{n,i}$, being the relative lever arm of $\mathrm{vP}_{i}$ to dot $n$. This stage can also be extended to integrate compensations for the cross-capacitance effects from the charge sensor to the QDs, if desired.
@@
with the entries $\alpha^{N}_{i,i}$ of the inverse matrix $M^{-1}_{3}$ defined as $d_{i}/V_{D}$ with $V_{D}$ the target charging voltage, and $d_{i}$ the charging voltage measured along the $\mathrm{O}_{i}$ voltage space.
@@
where the entry of $M^{-1}_{4}$ at the position $\alpha_{N_{\mathrm{B}}+i,j}^{\mathrm{OFF}}$ encodes the relative shift of the charge state of dot $i$ as the barrier $\mathrm{vB}_{j}$ voltage is changed.
Figure 1: Device architecture and virtualization workflow. (a) Layout of a ten-QD array based on a 3-4-3 geometry. Holes are trapped in gate-defined germanium QDs controlled by a set of barrier (dark blue), plunger (magenta), and screening (cyan) gates. (b) Schematic displaying the approximate potential landscape and position of the QDs in the array, with Dn, for $n\in[1,\dots,10]$, indicating each dot and SN, SE, SS, and SW indicating the north, east, south, and west charge sensor, respectively. (c) A typical 2D CSD of a double QD measured via the rf-reflectometry charge sensing on SN. (d) The workflow of the ML model: Each pixel is assigned a probability to be part of a horizontal, vertical, and interdot transition (diagonal and no transition classes not shown). (e) Probability distribution of each pixel to be a horizontal (left) or vertical (right) transition class. (f) Probability distribution of each pixel to be an interdot class. (g) A Gaussian fit to the log-transformed probability distribution for the interdot class. (h) Extracted coordinates of the interdot centers of mass based on dynamic thresholding.
The execution of fast two-qubit exchange-based gates requires modulating the tunnel coupling between adjacent QDs across several orders of magnitude [36, 22, 1]. A large voltage pulse on the barrier gates also affects the effective QD positions as the wave functions are brought close to overlap. This effect is expected to modify the aforementioned capacitive couplings limiting the range of validity of the virtual matrix extracted in the OFF-coupling regime. As a consequence, barrier control compensations dedicated to the high-coupling (ON) regime are introduced in layer 5. We anticipate that, depending on the targeted exchange interactions, linear compensations may not suffice and that nonlinear corrections to the plunger gates may be necessary when a barrier is pulsed substantially. Therefore, we heuristically define quadratically corrected virtual gates $\mathbf{K}$ as follows:
$\mathrm{K}_{j}=\mathrm{J}_{j}+\sum_{m}\left[\alpha^{\mathrm{ON}}_{j,m}{\mathrm{N}_{m}}^{2}+\beta^{\mathrm{ON}}_{j,m}\mathrm{N}_{m}\right],\vspace{-3pt}$ (5)
where the $(\sum_{m})$ involves only the nearest-neighboring plunger gates to barrier gate $j$, and $\alpha^{\mathrm{ON}}_{j,m}$, $\beta^{\mathrm{ON}}_{j,m}$ represent the first and second order compensation coefficients of barrier $J_{j}$ to the QDs $m$, respectively.
It is important to note that, in general, the charging energy of the QDs varies across different charge states. This means that the matrices and coefficients in Eqs. (1)–(5) are valid only locally, as full global virtualization across all charge states is not possible without many complex, nonlinear corrections. However, locally valid virtualization is sufficient for most applications since spin processors are operated in a predefined charge state. Thus, it is only necessary to perform virtualization after bringing the device to the desired regime of operation.
@@ II.2 Experimental setup
The full virtualization approach is demonstrated on a dense array of ten QDs defined in a planar germanium quantum well. This device [50] is based on low-disorder Ge/SiGe heterostructure grown on a Ge substrate, with the quantum well separated from the dielectric interface by a $55$-$\mathrm{nm}$ SiGe barrier [18].
The device uses a multilayer gate architecture to define an array of ten QDs arranged in a 3-4-3 configuration; see Fig. 1(a). Ten plunger gates and 12 barrier gates labeled $\mathrm{P}_{i}$ and $\mathrm{B}_{j}$ for $i\in[1,\dots,10]$ and $j\in[1,\dots,12]$, respectively, offer control of the array’s electrostatic potential landscape. Four additional plunger gates, labeled $\mathrm{P}_{N}$, $\mathrm{P}_{E}$, $\mathrm{P}_{W}$, and $\mathrm{P}_{S}$, control the sensor QD’s potentials. The ten QDs are labeled Dn, with $n\in[1,\dots,10]$, and their four peripheral sensor QDs are labeled SN, SE, SW, and SS, based on their cardinal directions; see Fig. 1(b). The additional eight screening gates screen the electric field from the plunger gates to prevent the formation of spurious QDs. The screening gates are omitted from the virtualization stack as they are not changed during normal device operation.
Prior to executing the virtualization stack, the QD array is pretuned to the few-hole regime, with each QD being either in a single-, triple-, or quintuple-hole occupancy for spin qubit manipulation [51]. The four charge sensors permit fast simultaneous radio-frequency- (rf)-reflectometry charge sensing in combination with video-mode acquisition and frequency multiplexing [52, 53].